If it takes 5 hours to go 100 miles, how long does it take to go 1 mile ?
Answer by lenny460(1073)

If it takes 5 hours to go 100 miles, how long does it take to go 1 mile ?
1 hour = 60 minutes
5 hours = 5 * 60 = 300 minutes
Therefore it takes 300 minutes to travel 100 miles
300 = 100
Divide each side by 100
300/100 = 100/100
3 minutes = 1 mile
Answer:
3 minutes



Lennox Obuong
Algebra Tutor